OpenStudy (anonymous):

multiply and simplify q^7 x q xq^-9x q-3 someone help :/

OpenStudy (solomonzelman):

It is not clear, can you rewrite it please?

OpenStudy (anonymous):

q^7 * q * q^-9 * q-3 ?

OpenStudy (ranga):

Is it \[\Large q ^{7} \times q \times q ^{-9} \times q ^{-3} \] ?

OpenStudy (solomonzelman):